# Towards Open Vocabulary Learning: A Survey

### v1

@@ 4.3 Open Vocabulary Instance Segmentation
Evaluation Metrics We adopt mean Average Precision (mAP) as the evaluation metric. For the constrained setting, we report base mAP and novel mAP separately. For the generalized setting, we report mAP for base, novel, and all classes.
Results on COCO. Tab. VII shows the open vocabulary instance segmentation results on the COCO dataset. The CGG method [38] achieves the best results on both constrained and generalized settings while not using pre-trained VLMs or any extra data. And Mask-free OVIS [214] gets a relatively high score on novel classes without any mask labels.
4.4 Open Vocabulary Panoptic Segmentation
@@
Evaluation Metrics Following previous works [38, 208, 175], we mainly adopt Panoptic Quality (PQ), Segmentation Quality (SQ), and Recognition Quality (RQ) as the evaluation metrics.
Results. Tab. VIII shows the results on the ADE20K dataset. ODISE-cap [175] achieves the best PQ score of 23.4. It surpasses the second-best score by 0.8. Tab. IX shows the results on the COCO dataset. PADing  [32] achieves a better PQ of 41.5 for seen classes, while Freeseg [216] achieves the highest PQ score of 29.8 for unseen classes.
5 CHALLENGES AND OUTLOOK
This section provides a short discussion for future research directions that can facilitate and envision the development of open vocabulary learning.
Explore Foundation Models With Specific Adapter For Custom Tasks. Vision foundation models [33, 16] can achieve good zero-shot performance on several standard classification and segmentation datasets. However, for several custom tasks, such as medical image analysis, and aerial images, there are still many corner cases. Thus, designing a task-specific adapter is needed for these custom tasks. Such adapters can fully utilize the knowledge of pre-trained foundation models.
Efficient Training For Target Dataset. As shown in the benchmark section 4, most state-of-the-art methods need huge data for pre-training to achieve good performances. However, the costs are expensive and even unavailable for many research groups to follow. Thus, with the aid of VLM, designing more efficient data learning pipelines or learning methods is more practical and affordable. One possible solution is to adopt the in-context learning [266, 267] to fully explore or connect the knowledge of VLMs and LLMs.
Base Classes Over-fitting Issues. Most approaches detect and segment novel objects by learning proposals from base class annotations. Thus, there are natural gaps in shapes and semantic information between novel objects and base objects. VLM models can bridge such gaps via pre-trained visual-text knowledge. However, most detectors still easily overfit the base classes when both the novel classes have similar shapes and semantics, since these classes are trained with higher confidence scores. More fine-grained feature discriminative modeling [268, 269, 270], including parts or attributes, is required to handle these problems.
6 Conclusion
This survey offers a detailed examination of the latest developments in open vocabulary learning in computer vision, which appears to be a first of its kind. We provide an overview of the necessary background knowledge, which includes fundamental concepts and introductory knowledge of detection, segmentation, and vision language pre-training. Following that, we summarize more than 50 different models used for various scene understanding tasks. For each task, we categorize the methods based on their technical viewpoint. Additionally, we provide information regarding several closely related domains. In the experiment section, we provide a detailed description of the settings and results. Finally, we suggest several future research directions for open vocabulary learning.

### v4

@@ A.3 Open Vocabulary Instance Segmentation
Evaluation Metrics We adopt mask mean Average Precision (mAP) as the evaluation metric. For the constrained setting, we report $AP_{base}$ and $AP_{novel}$ separately. For the generalized setting, we report mAP for the base, novel, and all classes.
Results on COCO. Tab. XIV shows the open vocabulary instance segmentation results on the COCO dataset. The CGG method [47] achieves the best results on both constrained and generalized settings while not using pre-trained VLMs or any extra data. And Mask-free OVIS [127] gets a relatively high score on novel classes without any mask labels.
A.4 Open Vocabulary Panoptic Segmentation
@@
Evaluation Metrics Following previous works [47, 86, 256, 133], we mainly adopt Panoptic Quality (PQ), Segmentation Quality (SQ), and Recognition Quality (RQ) as the evaluation metrics.
Results. Tab. XV shows the results on the ADE20K dataset. ODISE-cap [133] achieves the best PQ score of 23.4. It surpasses the second-best score by 0.8. Tab. XVI shows the results on the COCO dataset. PADing  [86] achieves a better PQ of 41.5 for seen classes, while Freeseg [131] achieves the highest PQ score of 29.8 for unseen classes.
A.5 Open Vocabulary Video Recognition
Settings. In video classification, existing methods usually test the zero-shot capability on downstream datasets (e.g., UCF and HMDB) pre-trained on Kinetics-400 [280] datasets. In most of the existing methods such as [270, 287], the ground truth of the Kinetics-400 is used for training, while recent method MAXI [123] does not require the ground truth when training on Kinetics-400.
Results Comparison. The comparison results are in Tab.V. Among the existing methods, Open-VCLIP [284] performs best on all three downstream datasets. It is also worth noting that the MAXI [123] performs well even without annotations from the Kinetics-400.
A.6 Open Vocabulary Video Instance Segmentation
Settings. To test the performance of open vocabulary video instance segmentation methods, MindVLT [57] collects LV-VIS datasets inheriting the categories of LVIS [49] and splits the categories into 659 base categories (frequent and common) and 553 novel categories. In the evaluation protocol, the LV-VIS is not used for training but for evaluation only. The mean Average Precision (mAP) is reported for comparison. $mAP_{base}$ and $mAP_{novel}$ refer to mAP for base and novel categories, respectively.
Results Comparison. The comparison results are in Tab.VI. Among the methods, MindVLT [57] achieves the best results on both base and novel categories. MindVLT [57] shows stronger improvement in the novel categories.
A.7 Open Vocabulary 3D Recognition
Settings. For open vocabulary 3D recognition, the zero-shot classification results on three different scale datasets, including ModelNet40 [330], ScanObjectNN [331], and Objaverse-LVIS [329] are reported. The first two datasets have 40 and 15 common categories, and the Objaverse-LVIS has 1156 LVIS [49] categories. The methods are trained on different datasets but tested on the three datasets for fair comparison.
Results Comparison. As in Tab.VII, using the large-scale OpenShape [276] datasets and larger backbones can significantly boost the performance on the downstream testing datasets.
A.8 Open Vocabulary 3D semantic segmentation
[…]
Results Comparison. As in Tab. VIII, RegionPLC [297] achieves remarkable performance on novel categories and thus indicates that it has a good capability to generalize to unseen categories.